OpenStudy (anonymous):

What percentage of a radioactive species would be found as daughter material after five half-lives? Please help me!!!

OpenStudy (farcher):

Every half life half of the parent nuclei decay and half are left. In the nest half life half of that half of parent nuclei decay.. Hence after 5 half lives the fraction of parent nuclei are left is \[\left(\frac 1 2 \right)^5\]so the fraction of daughter atoms which have been produced is \[1-\left ( \frac 1 2 \right )^5\]. So all you need to do now is convert that fraction into a percentage.